How many grams of potassium chloride are produced if 15 grams of potassium chlorate decomposes?"


1
Expert's answer
2021-06-07T03:13:08-0400

"2KClO_3\rightarrow 2KCl+3O_2"

By this equation we say

potassium chlorate and potassium chloride

Are in 1:1 ratio

Moles of KClO3="\frac{given \ mass}{molar \ mass}=\frac{15}{122.5}=0.12"

So mass of potassium chloride =0.12×74.5=8.94 g answer